An atomic nucleus initially moving at 390 m/s emits an alpha particle in the direction of its velocity, and the remaining nucleus slows to 250 m/s . Part A If the alpha particle has a mass of 4.0 u and the original nucleus has a mass of 222 u , what speed does the alpha particle have when it is emitted? Express your answer to two significant figures and include the appropriate units. v ′ α v α ′ = nothing nothing SubmitRequest Answer Provide Feedback Next
Answer #1

answer) by applying the conservation of ,momentum

total momentum of original nucleus m1=222*390 m/s=86580

now for remaining , the momentum is =250*218=54500

for alpha particle momentum is =4*v

so total m2=54500+4v

so since momentum is conserved

m1=m2

86580=54500+4v

4v=32080

v=8.0*103m/s

so answer is 8.0*103 m/s or 8.0 km/s or 8020 m/s